OpenStudy (anonymous):

Solve 2x-y=7 5x-4y=22

OpenStudy (anonymous):

\[x=2,y=-3\]

OpenStudy (anonymous):

\[8x-4y=28\] \[5x-4y=22\] Change the signs so \[-8x+4y=-28\]. Add the equations like a regular problem, canceling out the y and leaving -3x=-6, or x=2. Plug x back in the equation, and get 2(2)-y=7 or y=-3.

OpenStudy (anonymous):

Whoops, the first step should say multiply the top equation by 4 so the y cancels out.

OpenStudy (anonymous):

THANKS!